40g of a gas occupies 20 dm3 at 300K and 100 kilopascal pressure. If the pressure is changed to 50 kilopascal without changing the temperature, what would be its volume?

Grade:11

1 Answers

Vikas TU
7 years ago
That means the Temperature is constant.
Hence from ideal gas eqn. we get,
PV = nRT
100*20 = n*R*T
AND after changing pressure,
50*V = n*R*T
since both’s RHS is same.
PV = constant and hence  Boyle’s law is valid here.
Therefore,
100*20 = 50*V
V = 2000/50 => 40 dm3